Grade 11General Physics

A ball starts falling freely from a height h from a point A on an inclined plane with inclination alpha with the horizontal. After collision with the inclined plane, the ball rebounds elastically off the plane. Find the time after which it will strike the plane again and the distance from A.

Rituraj Singh
12 Years agoGrade 11

1 Answer

Arun Kumar
12 Years ago
Hello Student,
at the point v=root(2gh)
see the rebound velocity along the perpendicular will be same and opposite direction since elastic collision.
And horizontal momentum is conserved
so
we take x y acccording to the plane
vy=vcos\theta
vx=vsin\theta
ax=gsin\theta
ay=-gcos\theta
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty